Reports 1-1 of 1 Clear search Modify search
takafumi.ushiba - 15:40 Sunday 01 May 2022 (20631) Print this report
In-air health check of ITMY: check of the spectra of sensors

Spectra of each sensors were checked.
Though there is a problem of H2 geophone, which has been already known, the other sensors seem fine.

Figure 1 to 5 show the spectra of sensors on IP, BF, GAS, MN, IM stages, respectively.
H2 geophones on IP stage has a larger noise than the others but it is due to the ADC problem (klog21097 and its comments).
The other spectra, they seem fine compared to ETMY spectra.

We used spectra of REFNUM 0025 for this check.
Data were stored at /kagra/Dropbox/Measurements/VIS/SPECTRA/0/

Images attached to this report
Search Help